The diameters of ball bearings are distributed normally. The mean diameter is 74 millimeters and the standard deviation is 3 millimeters. Find the probability that the diameter of a selected bearing is greater than 71 millimeters. Round your answer to four decimal places.
Answer #1

Here, μ = 74, σ = 3 and x = 71. We need to compute P(X >= 71). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (71 - 74)/3 = -1

Therefore,
P(X >= 71) = P(z <= (71 - 74)/3)
= P(z >= -1)
= 1 - 0.1587
= 0.8413